OpenStudy (anonymous):

(1-cos)(1+cos)/cos^2 = tan^2

OpenStudy (anonymous):

what are you looking to do?

OpenStudy (anonymous):

finding a way to get that answer

OpenStudy (mertsj):

\[(1-\cos x)(1+\cos x)=1-\cos ^2x=\sin ^2x\]

OpenStudy (mertsj):

\[\frac{\cos ^2x}{\sin ^2x}=\tan ^2x\]